2.   a) Increases in real GDP are often interpreted as increases in the standard of living/welfare.     

         What are the main problems associated with this interpretation? (4 marks)

 In the Keynesian cross, assume that the consumption function is given by:

      C = 200 + 0.75 Yd

      Planned Investment is 100, Government purchases is 200 and Taxes are 100.

What does 200 represent? What is 0.75? what does Yd mean?

Compute the equilibrium level of income.

Calculate the multipliers of investment and taxes.

Compute autonomous spending. Assume that taxes are not payable when income is zero.

What level of Government spending is needed to achieve an income of 2,400?



1
Expert's answer
2020-12-31T09:30:39-0500

Solution:

2 a.). Real GDP as a measure of standard of living has a number of limitations which include the following:

·        Real GDP does not account for leisure time such as worker's vacation.

·        Real GDP does not include the actual levels of environmental cleanliness, health, and learning but rather what is spent on environmental protection, healthcare, and education.

·        Real GDP includes the cost of buying pollution-control equipment, but it does factor into account whether the air and water are actually cleaner or dirtier.

·        Real GDP includes spending on medical care, but it does not address whether life expectancy or infant mortality have risen or fallen.

·        Real GDP accounts for education spending, but it does not address directly how much of the population can read, write, or perform basic mathematics.


b.). 200 represents the Government spending.

      0.75 is the Marginal Propensity to Consume (MPC).

      Yd is disposable after-tax income, denoted as Y-T, where Y is the national income and T is tax.


c.). The equilibrium level of income:

At equilibrium: Y = AE

Y = C + I + G

Y = 200 + 0.75 (Y - T) + 100 + 200

Y = 200 + 0.75 (Y – 100) + 300

Y = 200 + 0.75Y – 75 + 300

Y = 200 + 300 – 75 + 0.75Y

Y = 425 + 0.75Y

Y – 0.75Y = 425

0.25Y = 425

Y = "\frac{425}{0.25}" = 1,700


d.). The Investment multiplier ="\frac{1}{1 - MPC}"

MPC = 0.75

"\frac{1}{1 - 0.75}" = "\frac{1}{0.25}" = 4


The Tax multiplier = "\frac{MPC}{1 - MPC}"

 = "\frac{0.75}{0.25}" = 3


e.). Autonomous spending:

C = a + by

C = 200 + 0.75Yd

= 200 + 0.75 (Y – 100)

= 200 + 0.75 (1700 – 100)

= 200 + 1275 – 75 = 1,400

C = 1,400

b = 0.75  Y = 1,700

Therefore:

C = a + by

1400 = a + (0.75*1700)

1400 = a + 1275

a = 1400 – 1275 = 125/=

Autonomous spending = 125/=

 

e.). The level of Government spending needed to achieve an income of 2,400 is:

Y = C + I + G

2,400 = 1,400 + 100 + G

2,400 = 1,500 + G

G = 2,400 – 1,500

G = 900/=

The level of Government spending needed to achieve an income of 2,400 = 900/=